# Damien Hutsemékers    Jean Manfroid    Emmanuël Jehin    Cyrielle Opitom    Michele Bannister    Juan Pablo Carvajal    Rosemary Dorsey    K Aravind    Baltasar Luco    Brian Murphy    Thomas H. Puzia    Rohan Rahatgaonkar

### v1

@@ 4 Discussion and conclusions
The expected blackbody equilibrium temperature of the cometary surface is approximately $T\simeq 280\;r_{h}^{-1/2}$ K, where $r_{h}$ is in au (Manfroid et al. 2021; Guzik & Drahus 2021; Puzia et al. 2025). At the distances at which the comets are observed, this temperature is far too low to vaporize silicate, sulfide, and metallic grains that contain NiI and FeI atoms. Therefore, the presence of NiI and FeI atoms in the cometary coma is extremely puzzling. Several hypotheses have been proposed to explain the presence of the NiI and FeI atoms at low surface temperatures as well as the high NiI/FeI abundance ratio, much higher that the solar system value. These scenarios were first presented by Manfroid et al. (2021), and discussed further by Bromley et al. (2021) and Rahatgaonkar et al. (2025). In particular, Manfroid et al. (2021) proposed that the NiI and FeI atoms could be released from Ni(CO)4 and Fe(CO)5 carbonyls, which are characterized by very low sublimation temperatures. The lower sublimation temperature of Ni(CO)4 with respect to Fe(CO)5 could also explain the high NiI/FeI ratios observed at low surface temperatures. The carbonyl scenario, although appealing, predicts a Ni/Fe ratio that strongly depends on the temperature and thus on the heliocentric distance. Unfortunately, such a dependence was not observed in the sample of solar system comets studied by Manfroid et al. (2021).
On the contrary, comet 3I exhibits a net variation in its NiI/FeI abundance ratio as a function of the heliocentric distance (Fig. 5). Fig. 8 shows the ratio of the sublimation rates of Ni(CO)4 and Fe(CO)5 computed as a function of $r_{h}$, together with the measured NiI/FeI abundance ratios. The sublimation rates were computed as a function of the temperature as done in Manfroid et al. (2021), and then as a function of $r_{h}$ using the relation $T=280\;r_{h}^{-1/2}$. Finally, the ratio of the sublimation rates was corrected for the solar system Ni/Fe abundance ratio, $\log_{10}$ (Ni/Fe) = 1.25-1.25 (Asplund et al. 2009). Bromley et al. (2021) showed that the photodissociation of Ni(CO)4 and Fe(CO)5 into NiI and FeI occurs at very similar rates for both carbonyls and at distances consistent with the observed spatial distributions of the atoms, making comparisons between atomic production rates and carbonyl sublimation rates meaningful. Remarkably, the carbonyl sublimation ratio reproduces the NiI/FeI ratio observed in several comets, but not all. In particular, it reproduces the ratio measured in comet 3I at the onset of FeI at $r_{h}\simeq$ 2.64 au. However, the rapid decrease of NiI/FeI with $r_{h}$ in comet 3I does not follow the theoretical curve (Fig. 8). The ratio of the carbonyl sublimation rates therefore appears as an upper limit to the observed NiI/FeI ratios. Rahatgaonkar et al. (2025) recorded the emergence of NiI, and found that its production rate changes with the heliocentric distance much faster than expected from purely radiative processes. As shown in Fig. 9, this steep dependence can be reproduced assuming that the NiI production rate is essentially controlled by the Ni(CO)4 carbonyl sublimation rate. The behavior of comet 3I suggests that carbonyl sublimation begins very early, well before the release of water and dust. At this stage, the surface temperature of the nucleus is close to the blackbody equilibrium temperature. As the comet approaches the Sun, more material is ejected, particularly dust, which can raise the temperature beyond the black-body equilibrium temperature, at least in localized regions (Prialnik et al. 2004; Groussin et al. 2013; Höfner et al. 2017). Since the carbonyl sublimation rate increases rapidly with the temperature, hot regions can dominate their sublimation, producing NiI and FeI atoms with a NiI/FeI abundance ratio lower than expected at the equilibrium temperature (see the dotted curve in Fig. 8). The absence of a net correlation between the NiI/FeI ratio and $r_{h}$ for solar system comets may simply be due to the dispersion of their temperatures and surface properties at a given heliocentric distance. If this interpretation is correct, the NiI/FeI ratio could be used to estimate the temperature of a cometary nucleus, or subregions at its surface.
@@
While this scenario provides a straightforward way to understand the behavior of NiI and FeI atoms in comets and supports the existence of carbonyls in cometary material, it is clear that there is more to the story, as the NiI/FeI ratio is also influenced by the carbon content, as demonstrated by the relationship between the NiI/FeI and C2/CN ratios (Fig. 8; see also Fig. 6). As the C2 depletion is primordial rather than evolutionary (Bair & Schleicher 2025), the NiI/FeI abundance ratio could also be a primordial property, at least in part. The possible link between the production of metal atoms and the abundance of carbon oxydes (Fig. 4), suggests that when more CO or CO2 molecules are available in the cometary material, more Ni and Fe could have been incorporated in carbonyls during the comet formation, so that more NiI and FeI atoms are finally produced when these carbonyls sublimate. A better understanding of the formation of carbonyls in various chemical environments would needed, in particular to see if an initial Ni(CO)4 / Fe(CO)5 asymmetry can already be produced during the comet formation, depending on the carbon content. Finally, it is worth to remember that other mechanisms, such as superheating of nanograins, could also contribute to the release of NiI and FeI atoms from refractory material, with a less extreme NiI/FeI ratio (see Manfroid et al. 2021). At the end, for Sun-grazing comets, complete vaporization of refractory grains occurs, leading to a solar NiI/FeI abundance ratio.
In conclusion, comet 3I, which is a C2-depleted comet, exhibits extreme properties in the early phases of its activity with regard to the production rates and abundance ratios of NiI and FeI. The rapid changes with the heliocentric distance suggest that comet 3I could soon become indistinguishable from solar system comets with respect to the NiI/FeI ratio. Nevertheless, comet 3I remains exceptional compared to solar system comets due to its high, asymmetrical production rate of NiI and FeI (Fig. 7). Compared to comets in the solar system, comet 3I is either metal-rich, which could potentially constrain the predictions of Hopkins et al. (2025), or, more likely, it has enhanced metal production due to chemical anomalies, such as the very high CO2/H2O abundance ratio (Cordiner et al. 2025).
@@ Acknowledgements.
DH and EJ are Research Directors at the F.R.S-FNRS. JM is honorary Research Director at the F.R.S-FNRS.
@@ Appendix A Tables
Table 2: FeI and NiI production rates, and their ratio. [^4] The data obtained on Sept. 3 and 4 in very similar circumstances were averaged to increase the signal-to-noise ratio.
Date nlines log10 Q(FeI) log10 Q(NiI) log10 [Q(NiI)/Q(FeI)]
yyyy-$m$-$d$ FeI / NiI s-1 s-1
2025-08-12 - / 20 << 21.83 22.82 $\pm$ 0.06 >> 0.99
2025-08-15 - / 14 << 22.14 22.74 $\pm$ 0.08 >> 0.60
2025-08-28 5 / 28 22.00 $\pm$ 0.08 23.27 $\pm$ 0.05 1.27 $\pm$ 0.10
2025-09-3/4 17 / 36 22.61 $\pm$ 0.04 23.51 $\pm$ 0.04 0.90 $\pm$ 0.06
2025-09-10 42 / 36 23.09 $\pm$ 0.03 23.75 $\pm$ 0.03 0.66 $\pm$ 0.04
2025-09-12 46 / 37 23.20 $\pm$ 0.02 23.80 $\pm$ 0.03 0.60 $\pm$ 0.04
Table 3: OH, CN, and C2 production rates. [^5] The data obtained on Sept. 3 and 4 in very similar circumstances were averaged to increase the signal-to-noise ratio.
Date log10 (OH) log10 Q(CN) log10 Q(C2)
yyyy-$m$-$d$ s-1 s-1 s-1
2025-08-12 - 24.0 $\pm$ 0.3 -
2025-08-15 - 24.35 $\pm$ 0.05 -
2025-08-28 26.37 $\pm$ 0.04 24.79 $\pm$ 0.04 24.6 $\pm$ 0.2
2025-09-3/4 26.59 $\pm$ 0.02 24.96 $\pm$ 0.03 24.6 $\pm$ 0.2
2025-09-10 27.05 $\pm$ 0.13 24.85 $\pm$ 0.01 24.36 $\pm$ 0.02
2025-09-11 - 25.16 $\pm$ 0.01 -
2025-09-12 27.15 $\pm$ 0.10 25.29 $\pm$ 0.01 24.47 $\pm$ 0.01
2025-09-14 - 25.26 $\pm$ 0.01 -

### v2

@@ 5 Discussion: The origin of NiI and FeI atoms
where $r_{h}$ is in astronomical units (Manfroid et al. 2021; Guzik & Drahus 2021; Puzia et al. 2025). At the distances at which the comets were observed, this temperature is far too low to vaporize silicate, sulfide, and metallic grains that contain NiI and FeI atoms. Therefore, the presence of NiI and FeI atoms in cometary comae is extremely puzzling. Several hypotheses have been proposed to explain the presence of the NiI and FeI atoms at low surface temperatures as well as the high NiI/FeI abundance ratio, much higher that the Solar System value. These scenarios were first presented by Manfroid et al. (2021), and discussed further by Bromley et al. (2021) and Rahatgaonkar et al. (2025). In particular, Manfroid et al. (2021) proposed that the NiI and FeI atoms could be released from Ni(CO)4 and Fe(CO)5 carbonyls, which are characterized by very low sublimation temperatures. The smaller sublimation temperature of Ni(CO)4 with respect to Fe(CO)5 could also explain the high NiI/FeI ratios observed at low surface temperatures. The carbonyl scenario, although appealing, predicts a NiI/FeI ratio that strongly depends on the temperature, and thus on the heliocentric distance. Unfortunately, such a dependence was not observed in the sample of Solar System comets studied by Manfroid et al. (2021).
On the contrary, comet 3I exhibits significant variation in its NiI/FeI abundance ratio as a function of the heliocentric distance, enabling us to further explore this scenario. Rahatgaonkar et al. (2025) already noticed that the production rate of NiI changes much faster with the heliocentric distance than can be explained by purely radiative processes.
Bromley et al. (2021) showed that the photodissociation of Ni(CO)4 and Fe(CO)5 into atomic NiI and FeI is very fast, with similar rates for both carbonyls, and occurs at nucleocentric distances consistent with the observed spatial distributions of the atoms. Therefore, the NiI and FeI production rates can be directly compared to the carbonyl sublimation rates. The sublimation rates of Ni(CO)4 and Fe(CO)5 were computed as a function of the temperature as in Manfroid et al. (2021). Then, they were expressed as a function of $r_{h}$ using Eq. 1. The ratio of the sublimation rates was finally corrected by assuming an intrinsic cometary Ni/Fe abundance ratio equal to the Solar System Ni/Fe abundance ratio [^5] It should be emphasized that the Ni/Fe abundance ratio in the cometary material (ices and dust) differs from the abundance ratio of the NiI and FeI atoms in the coma, which depends on the sublimation rates., $\log_{10}$ (Ni/Fe) = 1.25-1.25 (Asplund et al. 2009).
Figure 7 shows the variation in Q(NiI) between 3.78 and 1.85 au, and the variation in Q(FeI) between 2.64 au and 1.85 au. The solid curves represent the production rates of NiI and FeI computed from the sublimation rates of Ni(CO)4 and Fe(CO)5, which were multiplied by the effective emitting area of each species. The sublimation rates were computed assuming that the comet was at the equilibrium temperature at each heliocentric distance (Eq. 1). The effective emitting area was estimated empirically by shifting vertically (in logarithm) the sublimation rate curves (in molecules m-2 s-1) to match the observed production rates (in atoms s-1). The effective emitting areas were found to be equal to 1.6 m2 for both Ni(CO)4 and Fe(CO)5. This area is small, as is expected for minor species. However, its value should be interpreted with caution given the basic model. The Fe(CO)5 sublimation model nicely reproduces the variation in Q(FeI) with the heliocentric distance. For NiI, the global trend is reproduced, indicating that carbonyl sublimation begins very early. However, the increase in Q(NiI) with the heliocentric distance is less steep than the simple sublimation model predicts. This suggests that the temperature of the NiI-emitting regions may differ from the blackbody equilibrium temperature. Indeed, a reasonable fit can be obtained using the ad hoc relation $T=237\;r_{h}^{-1/3}$. As the comet approaches the Sun, more material is ejected, fractures form, and the temperature can deviate significantly from the blackbody equilibrium temperature, at least in localized regions (e.g., Prialnik et al. 2004; Groussin et al. 2013; Höfner et al. 2017). The different behavior of NiI and FeI with respect to the simple sublimation model could also indicate that these atoms are partially released from different subregions.
Figure 8 shows the ratio of the sublimation rates of Ni(CO)4 and Fe(CO)5 computed as a function of $r_{h}$ according to Eq. 1, together with the NiI/FeI abundance ratios measured in interstellar and Solar System comets. As before, the ratio of the sublimation rates is adjusted to account for the cometary intrinsic Ni/Fe abundance ratio, which is assumed to be equal to the Solar System value. Although the model is extremely basic, the computed ratio of the sublimation rates of Ni(CO)4 to Fe(CO)5 (solid curve) reproduces the NiI/FeI ratio observed in several comets within a factor 2-3, though not in all of them. In particular, it accurately reproduces the ratio observed in comet 3I at the onset of FeI at $r_{h}\simeq$ 2.64 au. However, the rapid decrease in the NiI/FeI ratio with $r_{h}$ does not follow the theoretical curve computed under the hypothesis that the comet remains at the equilibrium temperature. Assuming a different temperature dependence, such as $T=237\;r_{h}^{-1/3}$, allows us to reproduce the NiI/FeI variation (dotted curve). Therefore, the ratio of the carbonyl sublimation rates at the equilibrium temperature should be considered as an upper limit of the observed NiI/FeI ratios. The lack of correlation between the NiI/FeI ratio and $r_{h}$ for Solar System comets may be due to the dispersion of their temperatures and surface properties at a given heliocentric distance.
Figure 7: Q(NiI) and Q(FeI) production rates in comet 3I as a function of the heliocentric distance. Squares represent UVES measurements and triangles X-shooter measurements. The curves show the production rates of NiI (in red) and FeI (in blue), which were obtained by multiplying the sublimation rates of the Ni(CO)4 and Fe(CO)5 carbonyls by the effective emitting area of each species (see text). For the solid curves, the sublimation rates were computed assuming $T=280\;r_{h}^{-1/2}$. For the dotted red curve, the sublimation rate was computed assuming $T=237\;r_{h}^{-1/3}$ for Ni(CO)4.
Figure 8: Q(NiI)/Q(FeI) as a function of the heliocentric distance for interstellar and Solar System comets. The solid curve represents the ratio of the sublimation rates of the Ni(CO)4 and Fe(CO)5 carbonyls computed assuming $T=280\;r_{h}^{-1/2}$ for both Ni(CO)4 and Fe(CO)5. The dotted line represents the sublimation rate ratio assuming $T=237\;r_{h}^{-1/3}$ for the Ni(CO)4 emitting region, as in Fig. 7.
@@
Although incomplete, this scenario provides a straightforward explanation for the release of NiI and FeI atoms in comets and suggests the presence of carbonyls in the cometary material. However, it is clear that there is more to the story. The NiI/FeI ratio is also connected to the carbon content, as is shown by the relationship between the NiI/FeI and C2/CN ratios (Fig. 6). Since the C2 depletion is primordial rather than evolutionary (Bair & Schleicher 2025), the NiI/FeI abundance ratio could also be primordial in part. Comet 3I exhibits a CO2/H2O abundance ratio of 7.6 at $r_{h}$ = 3.3 au, which is higher than that of Solar System comets, and a CO/H2O ratio of 1.7, which is consistent with other cometary observations (Cordiner et al. 2025), including with those of comet 2I, which exhibits CO/H2O ratios ranging from 1.3 to 1.6 at 2.0 au (Bodewits et al. 2020). Comet C/2016 R2 exhibits more extreme abundance ratios: CO2/H2O $\simeq$ 30 and CO/H2O $\simeq$ 300 at $r_{h}$ = 2.8 au (McKay et al. 2019). The high Q(FeI+NiI) production rates observed in comets 3I and C/2016 R2 at approximately 2-3 au (Sect. 4 and Fig. 5) could be related to the high abundance of CO2, CO, or both. As is suggested by Cordiner et al. (2025), the high CO2/H2O measured in comet 3I could be due to a lower temperature of the cometary surface, which would also affect the metal production. On the other hand, when more CO2 and CO molecules are available in the cometary material, more NiI and FeI atoms could be incorporated into carbonyls during the comet formation, so that more NiI and FeI atoms are finally produced when these carbonyls sublimate. A better understanding of carbonyl formation in various chemical environments is needed, in particular to determine whether an initial Ni(CO)4 / Fe(CO)5 asymmetry can already be produced during the comet formation, depending on the carbon oxide content. Interestingly, Q(FeI+NiI) is dominated by NiI in comet 3I (Fig. 5), which has a high CO2/CO ratio, while Q(FeI+NiI) is dominated by FeI in comet C/2016 R2, which, on the contrary, has a high CO/CO2 ratio. Finally, it is worth noting that other mechanisms, such as superheating of nanograins, could also contribute to the release of NiI and FeI atoms from refractory material. However such mechanisms cannot fully explain the observed NiI/FeI ratios (see Manfroid et al. 2021). For Sun-grazing comets, complete vaporization of refractory grains occurs, leading to a solar NiI/FeI abundance ratio.
6 Conclusions
Since the beginning of its activity, the interstellar comet 3I/ATLAS has been regularly observed with UVES and X-shooter. These observations have provided an unprecedented dataset revealing the evolution of the NiI and FeI emission lines with heliocentric distance. During the initial stages of its activity, comet 3I exhibited extreme and unusual NiI/FeI abundance ratios. However, as its heliocentric distance decreased, the ratio became indistinguishable from those observed in Solar System comets and in the interstellar comet 2I/Borisov. Comet 3I was found to be C2-depleted, with a NiI/FeI abundance ratio finally consistent with other C2-depleted comets. Nevertheless, comet 3I remains exceptional due to its high total production rate of NiI and FeI, which is at least one order of magnitude larger than that of other comets. This enhanced metal production is possibly related to other chemical anomalies, such as the high CO2/H2O abundance ratio.
The NiI and FeI production rates, as well as their variations with heliocentric distance, are interpreted within the context of the sublimation of Ni(CO)4 and Fe(CO)5 carbonyls. This scenario provides a straightforward explanation for the asymmetric release of NiI and FeI atoms in the cometary coma and how it depends on heliocentric distance. Post-perihelion observations of NiI and FeI would be helpful to confirm the findings of this paper and clarify the trend with the heliocentric distance, in particular to discriminate between genuine variation with the heliocentric distance and effects due to processing of the outer layers of the nucleus.
@@ Acknowledgements.
We thank the referee for comments that helped significantly improve the manuscript. DH and EJ are Research Directors at the F.R.S-FNRS. JM is honorary Research Director at the F.R.S-FNRS. THP, RR, JPC, BL gratefully acknowledge support from the National Agency for Research and Development (ANID) under the following grants: CATA-Basal FB210003 and Beca de Doctorado Nacional (R.R. and J.P.C.). R.C.D. acknowledges support from grant #361233 awarded by the Research Council of Finland to M. Granvik.
@@ Appendix A Tables
Table 2: Observing circumstances: X-shooter. [^6] In Tables 1 and 2, $r_{h}$ and $\Delta$ are the heliocentric and geocentric distances of the comet. $\dot{r_{h}}$ and $\dot{\Delta}$ are the corresponding velocities. $w_{B}$, $w_{R}$, $h_{B}$, and $h_{R}$ refer to the blue/red slit width and height, respectively.
Date $r_{h}$ $\dot{r_{h}}$ $\Delta$ $\dot{\Delta}$ $w_{B}/h_{B}$
yyyy-$m$-$d$ au km s-1 au km s-1
2025-07-23 3.78 -56.8 2.96 -32.5 1.6 /11.0
2025-07-27 3.65 -56.6 2.89 -28.7 1.6 /11.0
2025-07-30 3.55 -56.4 2.84 -26.0 1.6 /11.0
2025-08-09 3.23 -55.6 2.72 -17.6 1.6 /11.0
2025-08-14 3.07 -55.0 2.67 -14.0 1.6 /11.0
2025-08-16 3.01 -54.8 2.65 -12.7 1.6 /11.0
2025-08-21 2.85 -54.1 2.62 -9.7 1.6 /11.0
2025-09-20 1.97 -45.4 2.53 -3.9 1.6 /11.0
2025-09-22 1.92 -44.3 2.52 -4.2 1.6 /11.0
2025-09-23 1.90 -43.7 2.52 -4.4 1.6 /11.0
2025-09-24 1.87 -43.1 2.52 -4.5 1.6 /11.0
2025-09-25 1.85 -42.5 2.52 -4.7 1.6 /11.0
Table 3: FeI and NiI production rates, and their ratio.
Date $r_{h}$ nlines log10 Q(FeI) log10 Q(NiI) log10 [Q(NiI)/Q(FeI)] Instrument
yyyy-$m$-$d$ au FeI / NiI s-1 s-1
2025-07-23 3.78 - / 2 - 22.11 $\pm$ 0.21 - X-shooter
2025-07-27 3.65 - / 2 - 22.35 $\pm$ 0.21 - X-shooter
2025-07-30 3.55 - / 4 - 22.43 $\pm$ 0.06 - X-shooter
2025-08-09 3.23 - / 3 - 22.83 $\pm$ 0.02 - X-shooter
2025-08-12 3.14 - / 20 << 21.83 22.82 $\pm$ 0.06 >> 0.99 UVES
2025-08-14 3.07 - / 11 - 22.83 $\pm$ 0.07 - X-shooter
2025-08-15 3.04 - / 14 << 22.14 22.74 $\pm$ 0.08 >> 0.60 UVES
2025-08-16 3.01 - / 12 - 22.91 $\pm$ 0.06 - X-shooter
2025-08-21 2.85 - / 14 - 23.08 $\pm$ 0.06 - X-shooter
2025-08-28 2.64 5 / 28 22.00 $\pm$ 0.08 23.27 $\pm$ 0.05 1.27 $\pm$ 0.10 UVES
2025-09-3/4 2.44 17 / 36 22.61 $\pm$ 0.04 23.51 $\pm$ 0.04 0.90 $\pm$ 0.06 UVES
2025-09-10 2.25 42 / 36 23.06 $\pm$ 0.03 23.75 $\pm$ 0.03 0.66 $\pm$ 0.04 UVES
2025-09-12 2.19 46 / 37 23.20 $\pm$ 0.02 23.80 $\pm$ 0.03 0.60 $\pm$ 0.04 UVES
2025-09-20 1.97 11 / 18 23.68 $\pm$ 0.03 24.05 $\pm$ 0.05 0.37 $\pm$ 0.06 X-shooter
2025-09-22 1.92 16 / 18 23.93 $\pm$ 0.03 24.25 $\pm$ 0.04 0.32 $\pm$ 0.05 X-shooter
2025-09-23 1.90 21 / 18 24.00 $\pm$ 0.02 24.30 $\pm$ 0.04 0.30 $\pm$ 0.05 X-shooter
2025-09-24 1.87 18 / 20 24.06 $\pm$ 0.03 24.29 $\pm$ 0.04 0.23 $\pm$ 0.05 X-shooter
2025-09-25 1.85 16 / 20 24.07 $\pm$ 0.03 24.31 $\pm$ 0.04 0.24 $\pm$ 0.05 X-shooter
Table 4: OH, CN, and C2 production rates. [^7] In Tables 3 and 4, the spectra obtained on September 3 and 4 in very similar circumstances were averaged to increase the signal-to-noise ratio. OH production rates were only measured in the UVES spectra. CN production rates before August 21, 2025 can be found in Rahatgaonkar et al. (2025).
Date $r_{h}$ log10 Q(OH) log10 Q(CN) log10 Q(C2) Instrument
yyyy-$m$-$d$ au s-1 s-1 s-1
2025-08-12 3.14 - 24.2 $\pm$ 0.3 - UVES
2025-08-15 3.04 - 24.59 $\pm$ 0.05 - UVES
2025-08-28 2.64 26.58 $\pm$ 0.04 24.79 $\pm$ 0.04 24.6 $\pm$ 0.2 UVES
2025-09-3/4 2.44 26.79 $\pm$ 0.02 24.96 $\pm$ 0.03 24.6 $\pm$ 0.2 UVES
2025-09-10 2.25 27.24 $\pm$ 0.13 25.03 $\pm$ 0.01 24.57 $\pm$ 0.02 UVES
2025-09-11 2.22 - 25.34 $\pm$ 0.01 - UVES
2025-09-12 2.19 27.33 $\pm$ 0.10 25.46 $\pm$ 0.01 24.68 $\pm$ 0.01 UVES
2025-09-14 2.14 - 25.43 $\pm$ 0.01 - UVES
2025-09-20 1.97 - 25.61 $\pm$ 0.02 - X-shooter
2025-09-22 1.92 - 25.70 $\pm$ 0.02 25.24 $\pm$ 0.05 X-shooter
2025-09-23 1.90 - 25.72 $\pm$ 0.02 25.32 $\pm$ 0.05 X-shooter
2025-09-24 1.87 - 25.73 $\pm$ 0.02 25.41 $\pm$ 0.05 X-shooter
2025-09-25 1.85 - 25.73 $\pm$ 0.02 25.39 $\pm$ 0.05 X-shooter